OpenStudy (decentnabeel):

Prove that :if (x+y)(y+z)(x+z)=0 x,y,z equal to not zero then 1/x+1/y+1/z=1/(x+y+z)

OpenStudy (decentnabeel):

@marcelie

OpenStudy (decentnabeel):

@zepdrix

OpenStudy (phi):

you can do it by cases. if (x+y)(y+z)(x+z)=0 and x,y,z \( \ne\) 0 then we have 3 cases: x+y=0 y+z=0 x+z=0

OpenStudy (phi):

or any combination of the three terms may be zero assume x+y=0 so x= -y then \[ \frac{1}{x}+\frac{1}{y}+\frac{1}{z} = - \frac{1}{y}+\frac{1}{y}+\frac{1}{z} \\ = \frac{1}{z} \] and because x+y=0 we can say \[ \frac{1}{z} = \frac{1}{x+y+z} \] and thus \[ \frac{1}{x}+\frac{1}{y}+\frac{1}{z} = \frac{1}{x+y+z} \] the other two cases give the same result (you could argue "by symmetry") to finish, show that if both (x+y) and (y+z)=0 you get the same result, (and similarly for any pair) and finally, if all 3 terms = 0

OpenStudy (phi):

though maybe we don't have to consider "multiple zeros" as the first proof should also cover those case.